Given a soda can with a volume of 36 and a diameter of 4, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank).


1
Expert's answer
2021-05-21T07:00:03-0400

"Solution: For ~this~ case ~what~ we~ should~ do~ is~ model~ the ~soda ~can ~as~ a ~cylinder.
\\We have then:
\\V=\pi r^2 h
\\Where,
\\r: radius ~ of ~the ~can
\\h: height ~of ~the ~can
\\From ~`here, we ~clear~ the ~value ~of ~the ~height:
\\Substituting ~values~ we~ have:
\\We~ are~ now~ looking~ for~ the~ volume ~of ~the ~cone.
\\We ~have~ then:
\\V=\frac{1}{3} \pi r^2 h
\\Substituting~values~we~have:
\\V=\frac{1}{3} \pi (2^2) (2.87)=12.02 \approx 12
\\Answer:The~volume ~of~ a ~cone~ that~ fits~ perfectly~ inside~ the~ soda~ can~ is ~V=12"